OpenStudy (anonymous):

express the result in standard form? 3i/4-i

myininaya (myininaya):

this one is all yours joe

OpenStudy (anonymous):

Lets multiply the top and bottom by the complex conjugate 4+i: \[\frac{3i}{4-i}*\frac{4+i}{4+i} = \frac{3i(4+i)}{4^{2}+1^{2}} = \frac{-3+12i}{17} = -\frac{3}{17}+\frac{12}{17}i\]

OpenStudy (anonymous):

12i/17-3/17